A plane is flying the 2553mile trip from los Angeles to Holohulu into a 60 mph headwind.if the speed of the plane in still air is 310mph.How far from Los Angeles is the plane's point of no retutn
Answer by mananth(16946)

given information
310 + 60 = 370 mph speed with the wind
310 - 60 - 250 mph speed against the wind
:
Let d = dist to LA from point of no return
Then
(2553-d) = dist to Honolulu from point of no return
:
: Time = dist/speed
:
Return to LA= Continue to Honolulu time
d/360 = 2553-d/250
Cross multiply:
250d = 360(2553-d)
:
250d = 919080 - 360d
:
610d = 919080
d = 1506.69
Check solution by finding the distance to LA and calculating the time to both:
2553-1506.69= = 1046.31 mi Honolulu
:
Time to L A: 1506.69/360 = 4.19 hrs
Time to Honolulu: 1046.31/250 = 4.19 hrs